OpenStudy (anonymous):

Let f(x)=5x^(2) Part 1 Compute the slope of the secant line joining the point of the graph of F whose x coordinates are x=-9 and x=-8.9. Why I got 89.5? I know somethings is wrong, Please help!!

jimthompson5910 (jim_thompson5910):

f(x)=5x^(2) f(-9)=5(-9)^(2) f(-9)=405

jimthompson5910 (jim_thompson5910):

f(x)=5x^(2) f(-8.9)=5(-8.9)^(2) f(-8.9)=396.05

jimthompson5910 (jim_thompson5910):

The slope of the secant line joining the point of the graph of F whose x coordinates are x=-9 and x=-8.9 is then... m = ( f(x2) - f(x1) )/(x2 - x1) m = ( f(-9) - f(-8.9) )/(-9 - (-8.9)) m = ( 405 - 396.05 )/(-9 - (-8.9)) m = -89.5

jimthompson5910 (jim_thompson5910):

So you lost a negative sign somewhere

OpenStudy (anonymous):

Yeah, I have to go back!!! I appreciate... Thank you!!

jimthompson5910 (jim_thompson5910):

np

OpenStudy (anonymous):

Quick question??

jimthompson5910 (jim_thompson5910):

what's that

OpenStudy (anonymous):

What will be the tangent of the graph if x=-9 It asking me to compute the slope i put =405 and it said is wrong

jimthompson5910 (jim_thompson5910):

find the slope of the secant line from x = -9 to x = -8.99 then find the slope of the secant line from x = -9 to x = -8.999 keep dragging that second point closer and closer to x = -9 As the second point gets closer to x = -9, the secant line slowly becomes the tangent line (to x = -9)

OpenStudy (anonymous):

Ok, so the answer can x=-9?

jimthompson5910 (jim_thompson5910):

no, the answer is not -9

OpenStudy (anonymous):

I have to graph it then.?

jimthompson5910 (jim_thompson5910):

no, you basically just need to compute the slopes of the secant lines above

jimthompson5910 (jim_thompson5910):

the slopes will slowly approach the slope of the tangent line at x = -9

OpenStudy (anonymous):

Now I am confuse, cause I got the other answer right! and it was the same as this one. :(

jimthompson5910 (jim_thompson5910):

can you imagine a secant line going through x = -9 and x = -8.9 ?

OpenStudy (anonymous):

No, I looking for more info trying to understand.. please help!

jimthompson5910 (jim_thompson5910):

Ok say we have the graph (this is just some arbitrary graph) 

jimthompson5910 (jim_thompson5910):

Now say we have the point A on the graph